# Dynamical similarity analysis uniquely captures how computations develop in RNNs

### v2

@@ 4.2 Testing ordinal response of metrics to compositional learning setup
Figure 3: DSA is the only metric capturing how representations develop during compositional learning. (a) RNN inputs and outputs in our setup. (b) Schematic of dynamical representations during ‘Anti task’. Different colors show different task conditions. (c) Same as (b) but for the ‘Delay task’. (d) Same as (c) but showing the compound (Master) ‘DelayAnti task’. (e) Schematic of observation in (Driscoll et al., 2024). (f) List of training conditions used in our test case. (g) We calculate the dissimilarity of all training groups to the ‘Master’ group with results shown by training schedule and metric. Color of boxplots refers to color used for training conditions in (f). Order at the bottom of the plot highlights the expected dissimilarity. Lines next to boxplots are standard errors.
@@
Metrics for describing representations are naturally most useful for cases where the amount of data to be analyzed is too large to allow for fully manual description, because of long time courses and model dimensionality. As such, we next want to move to a more complex test case that still allows us to make specific predictions about the normative relationships of representations across models. To construct this test case, we build on analyses of RNNs trained to learn compound tasks in a compositional fashion (Driscoll et al., 2024; Yang et al., 2019).
@@
In their work, Driscoll et al. (2024) train RNNs to learn simple tasks inspired by classical neuroscience experiments. For example, the ‘Delay task’ (Fig. 3a) requires networks to observe two streams of continual inputs (‘Stimulus Modality 1’ and ‘Stimulus Modality 2’) while the fixation signal is on, and then choose the stimulus with the higher average value once the fixation value disappears. There is a delay between the end of the stimulus period and the end of the fixation signal, during which the RNN needs to remember the response. Tasks like these are implemented in Neurogym (Molano-Mazon et al., 2022) and commonly used in computational neuroscience (details on tasks in Methods). The interesting fact for our purposes is that these simple tasks can easily be compositionally combined. For example, the ‘Delay task’ can be combined with an ‘Anti task’ which results in a ‘DelayAnti’ task where network do the same as in the ‘Delay task’ but the choice at the end of a trial is towards the weaker stimulus. Driscoll and colleagues observed that the representational dynamics within networks while solving a compound task is a combination of the dynamics of both separate tasks. So, the dynamics of ‘Delay’ (Fig. 3c) and ‘Anti’ (Fig. 3b) combine to ‘DelayAnti’ (Fig. 3d). They also observed that if they created a compound task based on subtasks, that the performance of networks when solving the compound task would increase with the number of subtasks they have been pretrained on (Fig. 3e). This case of gradually developing representations alongside gradually developing computations (i.e. task solving abilities) allows us to build a test case for representational metrics based on RNNs.
@@
Our test case is based on training a set of networks with different training schedule based on the setup by Driscoll et al. (2024). This is depicted in Fig. 3f. The baseline RNN is freely trained on a compound ‘Master task’ (called ‘M’ on figures) without any pretraining (top row Fig. 3f). This ‘Master’ network (red) can freely learn all dynamical motifs it needs to solve the task with its recurrent weights. We then also have a set of networks which are pretrained on partial sets of the subtasks and later trained on the Master task, but all their weights except for input weights are frozen between pretraining and training (‘Partial pretraining’ networks; green). These networks can only partially learn the needed dynamics in their recurrent weights and need to use their input weights for otherwise missing recurrent transformations. As stated before, Driscoll and colleagues observed that these networks only partially developed the dynamics of the Master network. We build additional special cases which are fully pretrained on all subtasks and are either frozen or not frozen after pretraining (purple and grey respectively). We also add networks which are only allowed to use their input weights to learn the Master task, but without any pretraining (‘Master frozen’; light blue). Lastly, we add a group of untrained networks which only observe the inputs with purely random weights (‘Untrained’; yellow). For each group of networks listed in Fig. 3f we train each of the 72 RNN with different hyperparameters (see Methods).
@@
Based on prior results (Driscoll et al., 2024) we can make normative predictions about how similar the representations of each network group should be to the representations in the ’Master’ group (Fig.3g, bottom). We would expect ‘Untrained’ to be the least similar, followed by ‘Partial pretraining’, followed by both fully pretrained groups (with no specific expected order between these two). Prior results do not allow for strong predictions about the ‘Master & frozen’ group. To test whether metrics comply with the ordinal predictions, we now measure the dissimilarity between the networks of each training group and their corresponding network in the ‘Master’ group. Note that prior results (Maheswaranathan et al., 2019) show that the specific network parameters such as activation functions can have strong influences on the specific dynamics, so that we compare similarities between networks with the same hyperparameters. The results of these analyses are shown in Fig. 3g for all metrics. We observe that only DSA shows the order we expect to see based on prior results. Calculating pairwise significance tests between all groups with the dissimilarity distribution of ‘Full Pretraining’ shows that all groups are significantly different from this baseline, except for ‘Full Pretraining & Unfrozen’ (all p-values FDR corrected; values given in Appendix 6.3 Table 1).
@@ 4.3 Testing metrics to identify task related computations alongside task related dynamics
We now want to use our RNN test case to see whether metrics directly link to the computational process of the network. The simplest way to summarize the differential behavior of two networks is two calculate their difference in task accuracy. Networks with more similar task accuracy likely implement more similar computations than networks with less similar task accuracy, especially in simple tasks such as ours where there are not multiple different strategies which allow for a correct solution. Hence, we want to test whether higher similarity in representations of two networks is predictive of a lower gap in accuracy. For this, we measure the dissimilarity between the ‘Master’ group and any other group as before. For each comparison of networks, we also capture their difference in task accuracy on a validation set of trials. Instead of just doing this at the end of training, we are now capturing the networks during six windows of training (10-25%, 25-40%, …, 85-100% of training based on epoch number; Fig. 4a) and comparisons of dissimilarity are done within these blocks of training. Within a group of training progression, we take the median measure of accuracy and dissimilarity as the measure of that group. This results in 3456 pairs of dissimilarity and accuracy values (72 network parameter setups, 8 training schedules, 6 comparisons across time, all multiplied) for each metric. We expect that a higher level of dissimilarity will be linked to a higher disparity in accuracy. Fig. 4b shows these relationships split by metric. We observe that, again, only DSA shows the expected pattern, with a significant positive relationship between these variables ($b=0.22,p=3.3\times 10^{-47},R^{2}=0.12$ ; other regressions in Appendix 6.4 Table 2). These analyses further strengthen the case for DSA. Not only does it behave as predicted to training schedules, but it also shows evidence of linking directly to the computations that networks are executing. Our test case here shows a capability of DSA that is difficult to highlight in test cases which only use simulated and non-task-solving dynamics.
@@ 4.4 Metrics’ responses to increasing overlap of the training schedule, across the duration of learning
With the new RNN test case we were able to show that DSA shows expected ordinal responses to training schedules and can link representational dynamics to computations. This tested expectations based on Driscoll et al. (2024). Using representational metrics carries the promise that one can precisely quantify the relationship between dynamics, instead of just relying on qualitative observation. In this section we want to demonstrate this by testing whether an increasing overlap in training schedule also causes an increasing alignment of representations. Additionally, we want to observe how this alignment of develops over the duration of training.
@@
We start with testing for the effect of a gradually overlapping training schedule on representations. To test this, we run a full set of pairwise comparisons between networks, meaning between each network from one specific training schedule is compare with networks of every other training schedule (Fig. 5a). For each network comparison, we quantify to which degree the two networks overlap in their training schedule. For example, a network pretrained on task A before being trained on compound task M, would share 50% with a network trained only on M and 66% with a network trained on A, C, and M. Fig. 5a highlight example comparisons. As before, we only ever compare networks with the same hyperparameters. For this analysis we only use networks that have completed training. The results depicted in Fig. 5c show that DSA can recognize the difference between full overlap of tasks, partial overlap of tasks, and no overlap of tasks, but does not seem to identify a gradual increase in overlap within the groups which have a partial overlap in tasks. CKA and Procrustes behave similarly (Appendix Fig. 8a and Fig. 8b). All regression parameters are in Appendix 6.5 Table 3.
@@
We additionally want to assess how representations develop over the course of learning. For this we use the prior analysis structure of comparing every training setup to the ‘Master’ setup over training time (Fig. 5b). We then plot dissimilarity values over training time, grouped by the pretraining ‘rank’, meaning how many of the subtasks making up the Master task were included in a network’s pretraining. Being not pretrained would be 0%, whereas being pretrained on B and C would make 66% (Fig. 5b). Fig. 5d shows the results for DSA. The 100% line (purple) shows that the reference model and the Master model become more similar as the Master learns the task, and hence is closer to the computations conducted by the pretrained model. This pattern is inverted for networks that have 0% task overlap (green): they start out similar when neither have learned to solve the task but become less similar as the master network learns. We again do not observe a gradual difference between different groups of partial overlapping training schedules. Other metrics are not reacting to this signal (Appendix Fig. 8c and Fig. 8d).
@@ 4.5 Using DSA and the established test case to analyze the learning process of state space models
To compare whether SSMs learn like RNNs, we apply the same test case we applied to RNNs in Fig. 3 to SSMs. As with RNNs, we use a wide set of network parameters (see Methods) and compare across training schedules (Fig. 3f) but within hyperparameters. We observe that Mamba learns quicker ($2.7$ times quicker than RNNs on the Master task) and models converge more reliably (99% convergence vs. 91% convergence on Master task). The results of the representational dissimilarity analyses are depicted in Fig. 6a, showing the same analysis as in RNNs (same plot as Fig. 3g). Mamba shows that the group that was trained on the Master task but with all weights frozen, except for the input weights, is the most different from representations observed in the Master network ($p=2.7\times 10^{-16}$; other p values in Appendix 6.6 Table 6). All other groups seem to roughly be in the same range of similarity to Master, though note ‘Partial Pretraining’ is also significantly more like the Master network then the ‘Full Pretraining’ is ($p=6.2\times 10^{-4}$). Regardless, ‘Master & Frozen’ clearly sticks out as the most different group. This means that all training groups produce roughly the same dynamics except for the group which is only allowed to learn the task through its input weights. This suggests that the very expressive Mamba architecture, when trained without freezing on a simple task such as our, does not learn by changing its hidden state dynamics but instead mostly learns through optimizing the transformation of the out-read layer (i.e. the last MLP layer). Given that the hidden state dynamics of most trained network look like the dynamics of the untrained network (p-values given in Appendix 6.6 Table 7) would lead to the prediction that the hidden state of Mamba during this simple task functions like the reservoir of a reservoir network (Lukoševičius & Jaeger, 2009). As before, this is a pattern only identified by DSA (see Appendix 6.6 Fig. 9 for CKA and Procrustes). Additional control analyses show that the RNN pattern from Fig. 3g holds true even when splitting the results into the RNN subtypes (Fig. 6b and Fig. 6c).
@@ 5 Discussion
Representational alignment metrics have become popular for comparing representations in neural systems across architectures and conditions. While they can be a helpful and scalable tool, recent work also has identified potential pitfalls when they react to spurious signals (Dujmović et al., 2023; Soni et al., 2024). Here we introduce two test cases which can serve as benchmarks for dynamical representation alignment metrics. Using these to compare DSA, CKA, and Procrustes reveals that DSA is the only metric that can identify how representation emerge throughout learning and how these then link to developing task solving abilities. Building on the combination of our test case and DSA also allows us to form a specific hypothesis about reservoir-like learning in Mamba models confronted with simple tasks.
@@
We put forward two specific test cases, where the simpler attractor-based test case is aimed at fundamental skills needed for the more complex test case. These cases naturally do not fully validate DSA and hence we shy away from calling our work a full benchmark. Instead, we see this as a first step into the direction of identifying a set of well-established empirical finding which can then form a full benchmark for dynamical representation metrics. In this process, the researchers could carefully choose what they want (or do not want) metrics to capture. We believe this step of gradually collecting test cases will be necessary to generate a better understanding of what metrics can and cannot do. There likely is going to be a trade-off between the complexity of test cases and the precision of normative predictions that can be made about representations, similar to what we observe in our two cases. We hope that our example can spark a productive back and forth between finding test cases and improving measures that will ultimately increase researchers’ confidence in the metrics.
@@
We also show that combinations of test cases with (partially) validated metrics can start to become important tools for studying architectural changes. Using our RNN-based test case and DSA, we can quickly generate two new observations: In RNNs we seem to find that the development of representations in networks as a function of partial pretraining does not seem to be gradual but instead blocked into relatively discrete groups of not-pretrained, partially pretrained, and fully-trained. Additionally, through applying our case to Mamba we observe results which seem to suggest that Mamba learns the relatively simple tasks of our test case in a reservoir-like way, meaning that it does not strongly change its already very expressive hidden state dynamics throughout learning and instead seems to use the layers following the hidden state for task related learning. This is unless we restrain the models so that they can only learn with their input weights, which then has a measurable effect in terms of changes to hidden state dynamics. Due to the exploratory character of these analyses, our results still need to be confirmed with more detailed analysis. If confirmed with further analysis, this would gradually increase our trust in these metrics.

### v3

@@ 4.2 Metrics’ ability to identify compositional dynamics during over training in RNNs
Figure 3: DSA, but not CKA or Procrustes, captures how representations develop during compositional learning. (a) RNN inputs and outputs in our setup. (b) Schematic of dynamical representations during ‘Anti task’. Different colors show different task conditions. (c) Same as (b) but for the ‘Delay task’. (d) Same as (c) but showing the compound (Master) ‘DelayAnti task’. (e) Schematic of observation in (Driscoll et al., 2024). (f) List of training conditions used in our test case. (g) We calculate the dissimilarity of all training groups to the ‘Master’ group with results shown by training schedule and metric. Color of boxplots refers to color used for training conditions in (f). Order at the bottom of the plot highlights the expected dissimilarity. Lines next to boxplots are standard errors.
@@
Metrics for describing representations are naturally most useful for cases where the amount of data to be analyzed is too large to allow for fully manual description, because of long time courses and model dimensionality. As such, we next want to move to a more complex test case that still allows us to make specific predictions about the expected relationships of representations across models. To construct this test case, we build on analyses of RNNs trained to learn simple cognitive tasks commonly used in computational neuroscience (Barbosa et al., 2023; Yang et al., 2019; Molano-Mazón et al., 2023; Proca et al., 2024; Aitken & Mihalas, 2023) and implemented in the Neurogym library (Molano-Mazon et al., 2022). More specifically, we build on results showing how RNNs learn compound tasks in a compositional fashion (Driscoll et al., 2024; Yang et al., 2019).
@@
In their work, Driscoll et al. (2024) show that when a task is constructed as a combination of subtasks, the developing overall representation similarly looks like a combination of the representations of these subtasks. As a specific example, the ‘Delay task’ (Fig. 3a) requires networks to observe two streams of continual inputs (‘Stimulus Modality 1’ and ‘Stimulus Modality 2’) while the fixation signal is on, and then choose the stimulus with the higher average value once the fixation value disappears. There is a delay between the end of the stimulus period and the end of the fixation signal, during which the RNN needs to remember the response. Driscoll et al. (2024) make use of the fact that these simple tasks can easily be combined. For example, the ‘Delay task’ can be combined with an ‘Anti task’ which results in a ‘DelayAnti’ task where network do the same as in the ‘Delay task’ but the choice at the end of a trial is towards the weaker stimulus. They observed that the representational dynamics within networks while solving a compound task is a combination of the dynamics of both separate tasks. So, the dynamics of ‘Delay’ (Fig. 3c) and ‘Anti’ (Fig. 3b) combine to ‘DelayAnti’ (Fig. 3d). To show this they used empirical bifurcation diagrams with fixed-point analyses. They also observed that if they created a compound task, the performance of networks when solving the compound task would increase with the number of subtasks they were pretrained on (Fig. 3e). Their results give us a very specific analytical understanding of how networks combine motifs to solve these tasks.
@@
How can we use this knowledge to test representational alignment metrics? The work outlined above makes the specific prediction about how the similarity of RNNs’ representations develops as an overlap of the training schedule. To test the metrics, we can replicate the training setup by Driscoll et al. (2024) and see whether metrics identify these similarities of representations. This is depicted in Fig. 3f. The baseline RNN is freely trained on a compound ‘Master task’ (called ‘M’ on figures) without any pretraining (top row Fig. 3f). This ‘Master’ network (red) can freely learn all dynamical motifs it needs to solve the task with its recurrent weights. We then also have a set of networks which are pretrained on partial sets of the subtasks and later trained on the Master task, but all their weights except for input weights are frozen between pretraining and training (‘Partial pretraining’ networks; green). These networks can only partially learn the needed dynamics in their recurrent weights and need to use their input weights for otherwise missing recurrent transformations. As stated before, Driscoll and colleagues observed that these networks only partially developed the dynamics of the Master network. We build additional special cases which are fully pretrained on all subtasks and are either frozen or not frozen after pretraining (purple and grey respectively). We also add networks which are only allowed to use their input weights to learn the Master task, but without any pretraining (‘Master frozen’; light blue). Lastly, we add a group of untrained networks which only observe the inputs with purely random weights (‘Untrained’; yellow). For each group of networks listed in Fig. 3f we train each of the 72 RNN with different hyperparameters (see Methods 6.1).
What are the specific predictions for how a metric should behave? Based on the prior results we can make ordinal predictions about how similar the representations of each network group should be to the representations in the ’Master’ group (Fig.3g, bottom). The ’Master’ group is here the reference group and represents the networks which can learn the compound task without constrains. As the computational structure of the tasks is known, we expect a higher alignment (with the ’Master’ group) of the conditions trained with the highest share of subtasks. We would expect ‘Untrained’ to be the least similar, followed by ‘Partial pretraining’, followed by both fully pretrained groups (with no specific expected order between these two). Prior results do not allow for strong predictions about the ‘Master & frozen’ group. If metrics cannot identify this order then this would lead us to conclude that they cannot uncover the computational structure of the tasks and thus not be able to correctly detect the development of representations in the networks.
@@
To test whether metrics match predictions, we measure the dissimilarity between the networks of each training group and their corresponding network in the ‘Master’ group. Note that prior results (Maheswaranathan et al., 2019) show that the network parameters such as activation functions can have strong influences on the dynamics, so that we compare similarities between networks with the same hyperparameters. The results of these analyses are shown in Fig. 3g for all metrics. We observe that only DSA shows the order we expect to see based on prior results. Calculating pairwise significance tests between all groups with the dissimilarity distribution of ‘Full Pretraining’ shows that all groups are significantly different from this baseline, except for ‘Full Pretraining & Unfrozen’ (all p-values FDR corrected; values in Appendix 6.3 Table 5). These results suggest that DSA and CKA do not identify meaningful differences between groups. This statement stays true even when comparing all the groups against each other and not only to ’Master’ (Appendix Fig 9).
@@ 4.3 Testing metrics to identify task related computations alongside task related dynamics
We now want to use our RNN test case to see whether metrics directly link to the computational process of the network. The simplest way to summarize the differential computations of two networks is to calculate their difference in task accuracy. Networks with more similar task accuracy likely implement more similar computations than networks with less similar task accuracy, especially in simple tasks such as ours where there are not multiple different strategies which allow for a correct solution. Hence, for a metric to capture networks’ computations, we expect a higher similarity in representations of two networks to be predictive of a lower gap in accuracy. To test this, we measure the dissimilarity between the ‘Master’ group and any other group as before. For each comparison of networks, we also capture their difference in task accuracy on a validation set of trials. Instead of just doing this at the end of training, we are now capturing the networks during six windows of training (10-25%, 25-40%, …, 85-100% of training based on epoch number; Fig. 4a) and comparisons of dissimilarity are done within these blocks of training. Within a group of training progression, we take the median measure of accuracy and dissimilarity as the measure of that group. This results in 3456 pairs of dissimilarity and accuracy values (72 network parameter setups, 8 training schedules, 6 comparisons across time, all multiplied) for each metric. We expect that a higher level of dissimilarity will be linked to a higher disparity in accuracy. Fig. 4b shows these relationships split by metric. We observe that, again, only DSA shows the expected pattern, with a significant positive relationship between these variables ($b=0.22,p=3.3\times 10^{-47},R^{2}=0.12$ ; other regressions in Appendix 6.4 Table 6). We do observe that even networks with the same accuracy value (i.e. $0\%$ accuracy difference) can still differ in their representations, which is likely caused by the difference in their training schedule. These analyses further strengthen the case for DSA. Not only does it behave as predicted to training schedules, but it also shows evidence of linking directly to the computations that networks are executing. Our test case here shows a capability of DSA that is difficult to highlight in test cases which only use simulated and non-task-solving dynamics.
@@ 4.4 Metrics’ responses to increasing overlap of the training schedule, across the duration of learning
With the new RNN test case we were able to show that DSA shows expected ordinal responses to training schedules and can link representational dynamics to computations. This tested expectations based on Driscoll et al. (2024). Using representational metrics carries the promise that one can precisely quantify the relationship between dynamics, instead of just relying on qualitative observation. In this section we want to demonstrate this by testing whether an increasing overlap in training schedule also causes an increasing alignment of representations. Additionally, we want to observe how this alignment develops over training. The following analysis are exploratory and that analyses by Driscoll et al. (2024) do not make specific predictions for these cases.
@@
We start with testing for the effect of a gradually overlapping training schedule on representations. To test this, we run a full set of pairwise comparisons between networks, meaning each network from one specific training schedule is compared with networks of every other training schedule (Fig. 5a). For each network comparison, we quantify to which degree the two networks overlap in their training schedule. For example, a network pretrained on task A before being trained on compound task M, would share 50% with a network trained only on M and 66% with a network trained on A, C, and M. Fig. 5a highlight example comparisons. As before, we only ever compare networks with the same hyperparameters. For this analysis we only use networks that have completed training. The results depicted in Fig. 5c show that DSA can recognize the difference between full overlap of tasks, partial overlap of tasks, and no overlap of tasks, but does not seem to identify a gradual increase in overlap within the groups which have a partial overlap in tasks. CKA and Procrustes behave similarly (Appendix Fig. 9a and Fig. 9b). All regression parameters are in Appendix 6.5 Table 7.
@@
We additionally want to assess how representations develop over the course of learning. For this we use the prior analysis of comparing every training setup to the ‘Master’ setup over training time (Fig. 5b). We then plot dissimilarity values over training time, grouped by the pretraining ‘rank’, meaning how many of the subtasks making up the Master task were included in a network’s pretraining. Being not pretrained would be 0%, whereas being pretrained on B and C would make 66% (Fig. 5b). In this analysis we use all possible pretrained model configurations and so we use the blue, green and purple groups from Fig 3f . Fig. 5d shows the results for DSA. The 100% line shows that the reference model and the Master model become more similar as the Master learns the task, and hence is closer to the computations conducted by the pretrained model than the ’Partial Pretraining’ group. This pattern is inverted for networks that have 0% task overlap (green): they start out similar when neither have learned to solve the task but become less similar as the master network learns. We again do not observe a gradual difference between different groups of partial overlapping training schedules. Other metrics are not reacting to this signal (Appendix Fig. 9).
@@ 4.5 Using DSA and the established test case to analyze the learning process of state space models
To compare whether SSMs learn like RNNs, we apply the same test case we applied to RNNs in Fig. 3 to SSMs. As with RNNs, we use a wide set of network parameters (see Methods) and compare across training schedules (Fig. 3f) but within hyperparameters. We observe that Mamba learns quicker ($2.7$ times quicker than RNNs on the Master task) and models converge more reliably (99% convergence vs. 91% convergence on Master task). The results of the representational dissimilarity analyses are depicted in Fig. 6a, showing the same analysis as in RNNs (same plot as Fig. 3g). Mamba shows that the group that was trained on the Master task but with all weights frozen, except for the input weights, is the most different from representations observed in the Master network ($p=2.7\times 10^{-16}$; other p values in Appendix 6.6 Table 10). All other groups seem to roughly be in the same range of similarity to Master, though note ‘Partial Pretraining’ is also significantly more like the Master network then the ‘Full Pretraining’ is ($p=6.2\times 10^{-4}$). Regardless, ‘Master & Frozen’ clearly sticks out as the most different group. This means that all training groups produce roughly the same dynamics except for the group which is only allowed to learn the task through its input weights. This suggests that the very expressive Mamba architecture, when trained without freezing on a simple task such as ours, does not learn by changing its hidden state dynamics but instead mostly learns through optimizing the transformation of the out-read layer (i.e. the last MLP layer). Given that the hidden state dynamics of most trained network look like the dynamics of the untrained network (p-values given in Appendix 6.6 Table 11) would lead to the prediction that the hidden state of Mamba during this simple task functions like the reservoir of a reservoir network (Lukoševičius & Jaeger, 2009). As before, this is a pattern only identified by DSA (see Appendix 6.6 Fig. 10 for CKA and Procrustes). Additional control analyses show that the RNN pattern from Fig. 3g holds true even when splitting the results into the RNN subtypes (Fig. 6b and Fig. 6c).
@@ 5 Discussion
Representational alignment metrics have become popular for comparing representations in neural systems across architectures and conditions. While they can be a helpful and scalable tool, recent work also has identified potential pitfalls when they react to spurious signals (Dujmović et al., 2023; Soni et al., 2024). Here we introduce two test cases which can serve as a first step towards a benchmark for dynamical representation alignment metrics. Using these to compare DSA, CKA, and Procrustes reveals that DSA is the only metric that can identify how compositional representations emerge throughout learning and how these are then linked to task behaviours. Building on the combination of our test case and DSA also allows us to form a specific hypothesis about reservoir-like learning in Mamba models confronted with simple tasks.
@@
We put forward two specific test cases, where the simpler attractor-based test case is aimed at fundamental skills needed for the more complex test case. These cases naturally do not fully validate DSA and hence we shy away from calling our work a full benchmark. Instead, we see this as a first step into the direction of identifying a set of well-established empirical finding which can then form a full benchmark for dynamical representation metrics. In this process, the researchers could carefully choose what they want (or do not want) metrics to capture. We believe this step of gradually collecting test cases will be necessary to generate a better understanding of what metrics can and cannot do. There likely is going to be a trade-off between the complexity of test cases and the precision of predictions that can be made about representations, similar to what we observe in our two cases. We hope that our example can spark a productive back and forth between finding test cases and improving measures that will ultimately increase researchers’ confidence in the metrics. As the set of test cases increases, we will also increasingly be sure that metrics will generalize to new and unseen cases. Until then, the risk of overfitting to our specific cases seems low, as the quantitative predictions of the first case are already captured by all metrics and the qualitative predictions of the second case intrinsically do not allow for overfitting, as there is no specific metric to overfit to.
@@
We also show that combinations of test cases with (partially) validated metrics can become important tools for studying architectural changes. Using our RNN-based test case, we can generate two new observations: In RNNs we seem to find that the development of representations in networks as a function of partial pretraining does not seem to be gradual but instead blocked into relatively discrete groups of not-pretrained, partially pretrained, and fully-trained. Additionally, through applying our case to Mamba we observe results which suggest that Mamba learns the simple tasks of our test case in a reservoir-like way, meaning that it does not strongly change its very expressive hidden state dynamics throughout learning and instead seems to use the layers following the hidden state for task related learning. This is unless we restrain the models so that they can only learn with their input weights, which then has a measurable effect in terms of changes to hidden state dynamics. Due to the exploratory character of these analyses, our results still need to be confirmed with more detailed analysis. If confirmed with further analysis, this would gradually increase our trust in these metrics.